STRUCTURE STUDY OF THE CHIRAL LACTIDE MOLECULES BY CHIRPED-PULSE FTMW SPEC-TROSCOPY

\underline{DANIEL P. ZALESKI}, JUSTIN L. NEILL, and BROOKS H. PATE, Department of Chemistry, University of Virginia, McCormick Rd., P.O. Box 400319, Charlottesville, VA 22904; EWA BIALKOWSKA-JAWORSKA and ZBIGNIEW KISIEL, Institute of Physics, Polish Academy of Sciences, Al. Lotnikw 32/46, 02-668 Warszawa, Poland.

Lactide is a six member cyclic diester with two chiral centers that forms from lactic acid in the presence of heat and an acid catalyst. It can form either a homo-chiral (RR) structure with both methyl groups equatorial or a hetero-chiral (RS) structure where one methyl group is equatorial and the other methyl group is axial. Structurally lactide is similar to lactic acid dimer; however, the kinked ring is covalently bonded and two waters are lost. And unlike lactic acid dimer, which has a very small dipole moment, the dipole moment of lactide is on the order of 3 Debye. Here the microwave spectra of the highly rigid homo- and hetero-chiral lactides are presented, which were first assigned in a heated lactic acid spectrum where the chemistry took place in the reservoir nozzles. Further isotopic information from a commercial sample of predominately homo-chiral lactide was obtained leading to a Kraitchman substitution structure of the homo-chiral lactide. Preliminary results of the cluster of homo-chiral lactide with one water molecule attached are also presented.